OpenStudy (anonymous):

8-(3+b)=b-9 step by step please?

OpenStudy (saifoo.khan):

8 - 2b - b = b - 9 DYS

OpenStudy (lgbasallote):

distribute the negative sign first 8-3-b = b-9 put all the b on one side 2b = 8-3+9 2b = 14 b =7

OpenStudy (anonymous):

:D Thanks!